# Finite-state transducers for substitution tilings

### v1

@@ 4.4 Putting it all together
Subject to Conjectures 1 and 2, the algorithms presented in this section combine into a general-purpose method for computing a deterministic finite-state transducer for tiling substitution systems.
@@ 5 Infinite supertile boundaries
An input sequence like this describes a tile $t$, and its output a neighbouring tile $t^{\prime}$, such that the two tiles have no supertile in common at all. The union of the patches of tiling deflated from $t$’s supertiles does not cover the whole plane, and the same is true for $t^{\prime}$. [Goodman-Strauss1999] describes these as ‘infinitely large supertiles’; one might also describe them as supertiles of infinite order. They are both, so we shall simply say ‘infinite supertile’.
@@
An example of an infinite supertile boundary occurs in the P2 Robinson triangle substitution system shown in Figure 1. From that figure we can see that the upper part of edge 1 of the $A$ triangle, under deflation, becomes edge 2 of a $B$ triangle. Conversely, the upper part of edge 2 of a $B$ becomes edge 1 of an $A$. So we can construct the following address, describing an $A$ tile whose sequence of supertile types alternates between $B$ and $A$ forever, and in particular describing edge 1 of the lowest-order $A$:
$\mathcal{A}=\bigl(A,1\bigr),\bigl(\xleftarrow{A}B\bigr),\bigl(\xleftarrow{B}A\bigr),\bigl(\xleftarrow{A}B\bigr),\bigl(\xleftarrow{B}A\bigr),\bigl(\xleftarrow{A}B\bigr),\bigl(\xleftarrow{B}A\bigr),\ldots$
@@
$\mathcal{B}=\bigl(B,2\bigr),\bigl(\xleftarrow{B}A\bigr),\bigl(\xleftarrow{A}B\bigr),\bigl(\xleftarrow{B}A\bigr),\bigl(\xleftarrow{A}B\bigr),\bigl(\xleftarrow{B}A\bigr),\bigl(\xleftarrow{A}B\bigr),\ldots$
@@
This is perhaps the simplest example of an infinite-order supertile boundary. It is also a good example of the principle that tile addresses of this kind tend to coincide with tilings of particular interest, because these tile addresses represent the two tiles at the centre of the famous ‘infinite cartwheel pattern’, an instance of P2 which not only has reflection symmetry but almost has 10-way rotation symmetry too [^6] This phenomenon of near-symmetry derives from the ‘singular pentagrids’ described in [DEBRUIJN198153]: the tiling is ultimately generated from a pentagrid which actually has the symmetry group of a decagon, but at the points where more than two pentagrid edges meet, gaps are created which can be filled with Penrose tiles in multiple ways, and no way to fill the gaps can preserve all the symmetry. We shall use the word ‘singular’ for other nearly-symmetric instances of substitution tilings, by analogy., in that most of the plane is invariant under a rotation of $36^{\circ}$, with only the centre of the image and a few thin spokes varying. Figure 15 shows this pattern, with the non-rotationally-symmetric parts shaded, and the line of true reflection symmetry running vertically down the centre.
@@ 5.1 Eventually periodic tile addresses
The address $\mathcal{A}$ shown above has the useful property of being eventually periodic. Apart from the initial special symbol containing an edge type instead of a supertile index, the rest of the address repeats the two symbols $\bigl(\xleftarrow{A}B\bigr),\bigl(\xleftarrow{B}A\bigr)$ for ever.
@@
For example, the address $\mathcal{A}$ in the previous section is part of the infinite supertile $\left[\bigl(\xleftarrow{B}A\bigr),\bigl(\xleftarrow{A}B\bigr)\right]^{*}$, whereas its neighbour $\mathcal{B}$ is part of the infinite supertile $\left[\bigl(\xleftarrow{A}B\bigr),\bigl(\xleftarrow{B}A\bigr)\right]^{*}$. The repeating sequence in each of these infinite supertiles is a cyclic rotation of the other, implying that the two supertiles can be obtained from one another by deflation. In this case they also appear in the same tiling of the whole plane with a boundary between them, but in other cases, they are entirely separate tilings.
@@ 6.3 Analyses of tilings
In particular, suppose we restrict our attention once again to the useful class of eventually periodic addresses. As mentioned in section 5.1, if an individual tile address is described as an initial segment and a repeating segment, then its containing infinite supertile is defined by the repeating segment alone. Pairs of infinite supertiles bordering on each other are defined by a cyclic path within the winnowed DFA. By searching for those cycles in increasing order of length, it is possible to automate the process of enumerating the infinite supertile boundaries in the eventually-periodic class. An analysis similar to this is begun for the Spectre tiling in [Paolini]; this DFA technique permits it to be automated for any tiling.
@@ 6.3.1 Finding out why a substitution system was ambiguous
To do this, we first make a DFA $\mathcal{D}$ which matches the regular language of tile addresses in the refined system, consisting of a start state, and a state for each (layer, tile type) pair, with transitions derived from the substitution system in an obvious manner. We aim to find a pair of distinct tile addresses $A,A^{\prime}$ accepted by this DFA which have the same image under the ‘forgetful’ map $\phi$ which discards tile subtypes to recover the corresponding address in the original unrefined system.
@@ 7.1 Penrose P2 and P3 tilings
Figures 19, 20 and 21 show the P2 triangle system in full; figures 25, 26 and 27 show the P2 whole-tile system; figures 22, 23 and 24 show the P3 triangle system; figures 28, 29 and 30 show the P3 whole-tile system.
@@ 7.2 Ammann-Beenker tiling
Figures 31, 32 and 33 show the Ammann-Beenker substitution system involving two rhombus tiles and the two triangular halves of the square, and Figure 34 shows its unambiguous refinement. Figures 35, 36 and 37 show the system with a whole square, and Figure 38 shows the unambiguous refinement of that system.
@@ 7.3 Spectre tiling
[Smith_Sep2024] presents two substitution systems for the Spectre tiling. One, shown in figure 2.1 of the paper, is a minimalist system containing two tile types: a single Spectre, and the ‘Mystic’ double tile consisting of two Spectres at $30^{\circ}$ to each other. The other, spread across figures 4.2, 5.1, and 5.3, involves nine hexagonal metatiles and eight different types of edge, with one set of deflation rules turning hexagons into more hexagons, and a second set turning them into Spectre tiles.
Both of these systems can be represented in the combinatorial form required as input to these algorithms. The 9-hexagon system is essentially in that form already, but the simpler Spectre/Mystic system is more difficult to represent, because the paper does not state how the edges of a single Spectre or Mystic correspond to the edges of the patch of tiles deflated from each one. However, such an edge mapping does exist; information equivalent to it is available in various sources such as [Cheritat], and the author is indebted to Pieter Mostert in particular for discussions involving the related ‘comet-rhomb tiling’ from which it can be derived.
@@
Figures 40 and 42 show the deflation diagrams for the two tiles in full [^8] As a practical matter, when generating a Spectre tiling from this system, one could add a base layer to the substitution system which separates the Mystic tile into its two Spectres. This is a simple transformation, and left to the reader.. The difficult spurs in the Mystic deflation are shown detached from the main tiling, with an indication of where they should properly start from.
The algorithms in this article take all of this complication in their stride. The adjacency maps for the two deflations describe the complicated spurs in the same way as described in section 2.2, mapping some $\mathbf{E}_{u,v}$ to each other, and whenever two pairs of edges coincide, making an arbitrary choice about which one to map to which [^9] The adjacency map in Figure 42 shows a particular choice of which edges to map to which, but this is not critical to the algorithm. For example, the diagram shows the mappings $\mathbf{E}_{15,-4}\leftrightarrow\mathbf{E}_{12,-2}$ and $\mathbf{I}_{2,9}\leftrightarrow\mathbf{I}_{4,4}$ on a pair of edges which coincide if the separate spur is replaced at its proper starting point (shown by a red diamond). The alternative choice $\mathbf{E}_{15,-4}\leftrightarrow\mathbf{I}_{4,4}$ and $\mathbf{I}_{2,9}\leftrightarrow\mathbf{E}_{12,-2}$, mapping the top side of each of those edges to the bottom side of the other one, would work equally well and make no difference to the success of the algorithms..
The alternative 9-hexagon system is not shown in full here, partly because [Smith_Sep2024] gives all the details required, and also because it is redundant, for the following reason.
@@
The unambiguous system generated by Algorithm 9 exactly corresponds to the 9-hexagon system: each of its 9 tile types corresponds to one of the hexagons, and every deflation diagram in the refined system matches the corresponding diagram in Figure 5.1 of [Smith_Sep2024], except that the metatile shapes are Spectres and Mystics instead of hexagons. Figure 16 shows these deflation diagrams in full, with the same tile names (and colours) as [Smith_Sep2024] uses for the corresponding hexagons.
@@
Running that algorithm reports that there is just one ambiguous infinite supertile, with three distinct addresses in the refined system which all look the same in the unrefined Spectre/Mystic system. In the unrefined system, the ‘simplest’ address of a tile in the ambiguous supertile is a Spectre tile which is subtile number 3 of a larger Spectre (in the arbitrary indexing of subtiles used in Figure 40), which is subtile 3 of its still larger parent, and so on for ever:
$\hbox{spectre}\xleftarrow{3}\hbox{spectre}\xleftarrow{3}\hbox{spectre}\xleftarrow{3}\hbox{spectre}\xleftarrow{3}\cdots$
@@
$\Psi\xleftarrow{3}\Psi\xleftarrow{3}\Psi\xleftarrow{3}\Psi\xleftarrow{3}\cdots$
$\Xi\xleftarrow{3}\Pi\xleftarrow{3}\Xi\xleftarrow{3}\Pi\xleftarrow{3}\cdots$
$\Pi\xleftarrow{3}\Xi\xleftarrow{3}\Pi\xleftarrow{3}\Xi\xleftarrow{3}\cdots$
@@
$\bigl(\Xi,0\bigr)\xleftarrow{3}\Pi\xleftarrow{3}\Xi\xleftarrow{3}\Pi\xleftarrow{3}\cdots$
$\bigl(\Pi,9\bigr)\xleftarrow{3}\Xi\xleftarrow{3}\Pi\xleftarrow{3}\Xi\xleftarrow{3}\cdots$
@@ 7.4 Hat tiling
[Smith_Jul2024] presents two substitution systems for the hat tiling. One, shown in figure 2.11 of the paper, is a minimalist system very similar to the Spectre/Mystic system, consisting of two tiles, of which one is a single hat, and the other is a fusion of a reflected hat with one of its neighbours, in exactly the same way that the Mystic is a fusion of the rare skew Spectre with a normally oriented neighbour. These deflate to patches of hats labelled $H_{8}$ and $H_{7}$ respectively. The other system consists of four metatiles $H,T,P,F$ in simple polygonal shapes, with one deflation rule to turn them into more metatiles, and another to turn them into hats (shown in figures 2.8 and 4.1 of the paper respectively).
@@
Transforming each edge of the two tile types via these rules expands each one into a patch which can be tiled by one double-hat fusion and either 6 or 7 ordinary hats. Figures 45 and 47 show the results of this process [^11] As with the Spectre system, one might wish to add a base layer to this system separating the double-hat tile back into its two individual hats. Another useful optional base layer, in this or the HTPF system, is one that divides each hat into the eight kites of the underlying kite tiling, which allows the output of these algorithms to be handled in a discrete manner as subsets of the kites in that grid..
@@
Edge annotations for the non-overlapping HTPF system are derived in [Baake_2025] (figures 4 and 5), and reproduced here in Figure 49 (showing the edges subdivided as mentioned in section 2.3, with edge types and directions), and Figures 50 and 51 showing the rules for transforming each edge type into a sequence of sub-edges during deflation of metatiles to other metatiles or hats respectively. The resulting adjacency maps are shown in Figures 52 and 53.
@@
The third copy of this infinite supertile, involving the $h_{3}$ address, is part of a second 6-way singular hat tiling. If the two tilings are overlaid so that the $h_{3}$ supertile corresponds to the $h_{4}$ supertile, then the two tilings differ only in a single straight line of hats; if instead the two tilings are overlaid so that the $h_{3}$ supertile corresponds to $h_{2}$, then they instead differ in a wiggly path. [Socolar_2023] exhibits these two tilings and their two minimal differing paths.
@@ 8 Future possibilities
Such a transducer would not necessarily be able to keep the invariant seen in the transducers here, of emitting exactly one output symbol per input symbol, because not all substitution systems for the same tiling deflate by the same factor. For example, there is a ‘half-deflation’ system for the P2 and P3 Penrose tilings [DEBRUIJN198139] which deflates each one to the other, such that two deflations return to the same tiling you started with, having performed what would normally be a full deflation. [mathBlock2025] also presents systems for the hat tiling which take two deflations to repeat the effect of a single deflation in the $\{H_{7},H_{8}\}$ or HTPF systems (although in this case the intermediate tiling is still a hat tiling).
If an algorithm of this form were possible, then it might also be possible to use it to match a substitution system against itself, to determine if there exist any pairs of distinct addresses which give rise to an identical tiling of the plane. For example, the system in [mathBlock2025] has such pairs, because in some cases, half-inflating a perfectly symmetric instance of the hat tiling leads to a ‘singular’ (nearly symmetric) instance, and so the tile addresses in the symmetric tiling must break its symmetry in order to specify which way round the not-quite-symmetric singular deflation will be.

### v2

@@ 4.4 Putting it all together
The algorithms presented in this section combine into a general-purpose method for computing a deterministic finite-state transducer for tiling substitution systems.
@@ 5 Infinite supertile boundaries
An input sequence like this describes a tile $t$, and its output a neighbouring tile $t^{\prime}$, such that the two tiles have no supertile in common at all. The union of the patches of tiling deflated from $t$’s supertiles does not cover the whole plane, and the same is true for $t^{\prime}$. [7] describes these as ‘infinitely large supertiles’; one might also describe them as supertiles of infinite order. They are both, so we shall simply say ‘infinite supertile’.
@@
An example of an infinite supertile boundary occurs in the P2 Robinson triangle substitution system shown in Figure 1. From that figure we can see that the upper part of edge 1 of the $A$ triangle, under deflation, becomes edge 2 of a $B$ triangle (which we have indexed as subtile number 1 in figure 2). Conversely, the upper part of edge 2 of a $B$ becomes edge 1 of an $A$ (again subtile number 1). So we can construct the following address, describing an $A$ tile whose sequence of supertile types alternates between $B$ and $A$ forever, and in particular describing edge 1 of the lowest-order $A$:
$\mathcal{A}=\bigl(A,1\bigr),\bigl(\xleftarrow{1}B\bigr),\bigl(\xleftarrow{1}A\bigr),\bigl(\xleftarrow{1}B\bigr),\bigl(\xleftarrow{1}A\bigr),\bigl(\xleftarrow{1}B\bigr),\bigl(\xleftarrow{1}A\bigr),\ldots$
@@
$\mathcal{B}=\bigl(B,2\bigr),\bigl(\xleftarrow{1}A\bigr),\bigl(\xleftarrow{1}B\bigr),\bigl(\xleftarrow{1}A\bigr),\bigl(\xleftarrow{1}B\bigr),\bigl(\xleftarrow{1}A\bigr),\bigl(\xleftarrow{1}B\bigr),\ldots$
@@
This is perhaps the simplest example of an infinite-order supertile boundary. It is also a good example of the principle that tile addresses of this kind tend to coincide with tilings of particular interest, because these tile addresses represent the two tiles at the centre of the famous ‘infinite cartwheel pattern’, an instance of P2 which not only has reflection symmetry but almost has 10-way rotation symmetry too [^8] This phenomenon of near-symmetry derives from the ‘singular pentagrids’ described in [6]: the tiling is ultimately generated from a pentagrid which actually has the symmetry group of a decagon, but at the points where more than two pentagrid edges meet, gaps are created which can be filled with Penrose tiles in multiple ways, and no way to fill the gaps can preserve all the symmetry. We shall use the word ‘singular’ for other nearly-symmetric instances of substitution tilings, by analogy., in that most of the plane is invariant under a rotation of $36^{\circ}$, with only the centre of the image and a few thin spokes varying. Figure 16 shows this pattern, with the non-rotationally-symmetric parts shaded, and the line of true reflection symmetry running vertically down the centre.
The example transducer in Figure 8 demonstrates that these two addresses $\mathcal{A},\mathcal{B}$ are indeed mapped to one another, without any difficulty arising from the infinite supertile boundary: the symbols of address $\mathcal{B}$ shown above are generated as output from the transducer, if one follows transitions from start on the symbols of $\mathcal{A}$, and vice versa.
@@ 5.1 Eventually periodic tile addresses
The address $\mathcal{A}$ shown above has the useful property of being eventually periodic. Apart from the initial special symbol containing an edge type instead of a supertile index, the rest of the address repeats the two symbols $\bigl(\xleftarrow{1}B\bigr),\bigl(\xleftarrow{1}A\bigr)$ for ever.
@@
For example, the address $\mathcal{A}$ in the previous section is part of the infinite supertile $\left[\bigl(\xleftarrow{1}A\bigr),\bigl(\xleftarrow{1}B\bigr)\right]^{*}$, whereas its neighbour $\mathcal{B}$ is part of the infinite supertile $\left[\bigl(\xleftarrow{1}B\bigr),\bigl(\xleftarrow{1}A\bigr)\right]^{*}$. The repeating sequence in each of these infinite supertiles is a cyclic rotation of the other, implying that the two supertiles can be obtained from one another by deflation. In this case they also appear in the same tiling of the whole plane with a boundary between them, but in other cases, they are entirely separate tilings.
@@ 6.3 Analyses of tilings
In particular, suppose we restrict our attention once again to the useful class of eventually periodic addresses. As mentioned in section 5.1, if an individual tile address is described as an initial segment and a repeating segment, then its containing infinite supertile is defined by the repeating segment alone. Pairs of infinite supertiles bordering on each other are defined by a cyclic path within the winnowed DFA. By searching for those cycles in increasing order of length, it is possible to automate the process of enumerating the infinite supertile boundaries in the eventually-periodic class. An analysis similar to this is begun for the Spectre tiling in [9]; this DFA technique permits it to be automated for any tiling.
@@ 6.3.1 Finding out why a substitution system was ambiguous
To do this, we first make a DFA which matches the ‘diagonal’ subset $\mathcal{D}\subset\mathcal{T}\times\mathcal{T}$ of pairs of tile addresses in the refined system: the language of strings made by zipping the same tile address with itself. This is matched by a DFA isomorphic to the one for $\mathcal{T}$ itself, consisting of a start state, and a state for each $(\textit{layer},\textit{tile type})$ pair, with transitions derived from the substitution system in an obvious manner. We aim to find a pair of distinct tile addresses $A,A^{\prime}$ accepted by this DFA which have the same image under the ‘forgetful’ map $\phi$ which discards tile subtypes to recover the corresponding address in the original unrefined system.
@@ 7.1 Penrose P2 and P3 tilings
Figures 20, 21 and 22 show the P2 triangle system in full; figures 26, 27 and 28 show the P2 whole-tile system; figures 23, 24 and 25 show the P3 triangle system; figures 29, 30 and 31 show the P3 whole-tile system.
@@ 7.2 Ammann-Beenker tiling
Figures 32, 33 and 34 show the Ammann-Beenker substitution system involving two rhombus tiles and the two triangular halves of the square, and Figure 35 shows its unambiguous refinement. Figures 36, 37 and 38 show the system with a whole square, and Figure 39 shows the unambiguous refinement of that system.
@@ 7.3 Spectre tiling
[11] presents two substitution systems for the Spectre tiling. One, shown in figure 2.1 of the paper, is a minimalist system containing two tile types: a single Spectre, and the ‘Mystic’ double tile consisting of two Spectres at $30^{\circ}$ to each other. The other, spread across figures 4.2, 5.1, and 5.3, involves nine hexagonal metatiles and eight different types of edge, with one set of deflation rules turning hexagons into more hexagons, and a second set turning them into Spectre tiles.
Both of these systems can be represented in the combinatorial form required as input to these algorithms. The 9-hexagon system is essentially in that form already, but the simpler Spectre/Mystic system is more difficult to represent, because the paper does not state how the edges of a single Spectre or Mystic correspond to the edges of the patch of tiles deflated from each one. However, such an edge mapping does exist; information equivalent to it is available in various sources such as [1], and the author is indebted to Pieter Mostert in particular for discussions involving the related ‘comet-rhomb tiling’ from which it can be derived.
@@
Figures 41 and 43 show the deflation diagrams for the two tiles in full [^10] As a practical matter, when generating a Spectre tiling from this system, one could add a base layer to the substitution system which separates the Mystic tile into its two Spectres. This is a simple transformation, and left to the reader.. The difficult spurs in the Mystic deflation are shown detached from the main tiling, with an indication of where they should properly start from.
The algorithms in this article take all of this complication in their stride. The adjacency maps for the two deflations describe the complicated spurs in the same way as described in section 2.3, mapping some $\mathbf{E}_{u,v}$ to each other, and whenever two pairs of edges coincide, making an arbitrary choice about which one to map to which [^11] The adjacency map in Figure 43 shows a particular choice of which edges to map to which, but this is not critical to the algorithm. For example, the diagram shows the mappings $\mathbf{E}_{15,-4}\leftrightarrow\mathbf{E}_{12,-2}$ and $\mathbf{I}_{2,9}\leftrightarrow\mathbf{I}_{3,4}$ on a pair of edges which coincide if the separate spur is replaced at its proper starting point (shown by a red diamond). The alternative choice $\mathbf{E}_{15,-4}\leftrightarrow\mathbf{I}_{3,4}$ and $\mathbf{I}_{2,9}\leftrightarrow\mathbf{E}_{12,-2}$, mapping the top side of each of those edges to the bottom side of the other one, would work equally well and make no difference to the success of the algorithms..
The alternative 9-hexagon system is not shown in full here, partly because [11] gives all the details required, and also because it is redundant, for the following reason.
@@
The unambiguous system generated by Algorithm 9 exactly corresponds to the 9-hexagon system: each of its 9 tile types corresponds to one of the hexagons, and every deflation diagram in the refined system matches the corresponding diagram in Figure 5.1 of [11], except that the metatile shapes are Spectres and Mystics instead of hexagons. Figure 17 shows these deflation diagrams in full, with the same tile names (and colours) as [11] uses for the corresponding hexagons.
@@
Running that algorithm reports that there is just one ambiguous infinite supertile, with three distinct addresses in the refined system which all look the same in the unrefined Spectre/Mystic system. In the unrefined system, the ‘simplest’ address of a tile in the ambiguous supertile is a Spectre tile which is subtile 7 of a larger Spectre (in the arbitrary indexing of subtiles used in Figure 41), which is subtile 7 of its still larger parent, and so on for ever:
$\hbox{spectre}\xleftarrow{7}\hbox{spectre}\xleftarrow{7}\hbox{spectre}\xleftarrow{7}\hbox{spectre}\xleftarrow{7}\cdots$
@@
$\Psi\xleftarrow{7}\Psi\xleftarrow{7}\Psi\xleftarrow{7}\Psi\xleftarrow{7}\cdots$
$\Xi\xleftarrow{7}\Pi\xleftarrow{7}\Xi\xleftarrow{7}\Pi\xleftarrow{7}\cdots$
$\Pi\xleftarrow{7}\Xi\xleftarrow{7}\Pi\xleftarrow{7}\Xi\xleftarrow{7}\cdots$
@@
$\bigl(\Xi,0\bigr)\xleftarrow{7}\Pi\xleftarrow{7}\Xi\xleftarrow{7}\Pi\xleftarrow{7}\cdots$
$\bigl(\Pi,9\bigr)\xleftarrow{7}\Xi\xleftarrow{7}\Pi\xleftarrow{7}\Xi\xleftarrow{7}\cdots$
@@ 7.4 Hat tiling
[12] presents two substitution systems for the hat tiling. One, shown in figure 2.11 of the paper, is a minimalist system very similar to the Spectre/Mystic system, consisting of two tiles, of which one is a single hat, and the other is a fusion of a reflected hat with one of its neighbours, in exactly the same way that the Mystic is a fusion of the rare skew Spectre with a normally oriented neighbour. These deflate to patches of hats labelled $H_{8}$ and $H_{7}$ respectively. The other system consists of four metatiles $H,T,P,F$ in simple polygonal shapes, with one deflation rule to turn them into more metatiles, and another to turn them into hats (shown in figures 2.8 and 4.1 of the paper respectively).
@@
Transforming each edge of the two tile types via these rules expands each one into a patch which can be tiled by one double-hat fusion and either 6 or 7 ordinary hats. Figures 46 and 48 show the results of this process [^13] As with the Spectre system, one might wish to add a base layer to this system separating the double-hat tile back into its two individual hats. Another useful optional base layer, in this or the HTPF system, is one that divides each hat into the eight kites of the underlying kite tiling, which allows the output of these algorithms to be handled in a discrete manner as subsets of the kites in that grid..
@@
Edge annotations for the non-overlapping HTPF system are derived in [3] (figures 4 and 5), and reproduced here in Figure 50 (showing the edges subdivided as mentioned in section 2.4, with edge types and directions), and Figures 51 and 52 showing the rules for transforming each edge type into a sequence of sub-edges during deflation of metatiles to other metatiles or hats respectively. The resulting adjacency maps are shown in Figures 53 and 54.
@@
The third copy of this infinite supertile, involving the $h_{3}$ address, is part of a second 6-way singular hat tiling. If the two tilings are overlaid so that the $h_{3}$ supertile corresponds to the $h_{4}$ supertile, then the two tilings differ only in a single straight line of hats; if instead the two tilings are overlaid so that the $h_{3}$ supertile corresponds to $h_{2}$, then they instead differ in a wiggly path. [13] exhibits these two tilings and their two minimal differing paths.
@@ 8 Future possibilities
Such a transducer would not necessarily be able to keep the invariant seen in the transducers here, of emitting exactly one output symbol per input symbol, because not all substitution systems for the same tiling deflate by the same factor. For example, there is a ‘half-deflation’ system for the P2 and P3 Penrose tilings [5] which deflates each one to the other, such that two deflations return to the same tiling you started with, having performed what would normally be a full deflation. [14] also presents systems for the hat tiling which take two deflations to repeat the effect of a single deflation in the $\{H_{7},H_{8}\}$ or HTPF systems (although in this case the intermediate tiling is still a hat tiling).
If an algorithm of this form were possible, then it might also be possible to use it to match a substitution system against itself, to determine if there exist any pairs of distinct addresses which give rise to an identical tiling of the plane. For example, the system in [14] has such pairs, because in some cases, half-inflating a perfectly symmetric instance of the hat tiling leads to a ‘singular’ (nearly symmetric) instance, and so the tile addresses in the symmetric tiling must break its symmetry in order to specify which way round the not-quite-symmetric singular deflation will be.
9 Implementation
An implementation of the algorithms in this article, written in Rust, is in the crates.io package repository under the name substitution-tiling-transducers.